# Iterative assembly of ${}^{171}$b atom arrays in cavity-enhanced optical lattices

### v2

@@ Abstract
Assembling and maintaining large arrays of individually addressable atoms is a key requirement for continued scaling of neutral-atom-based quantum computers and simulators. In this work, we demonstrate a new paradigm for assembly of atomic arrays, based on a synergistic combination of optical tweezers and cavity-enhanced optical lattices, and the incremental filling of a target array from a repetitively filled reservoir. In this protocol, the tweezers provide microscopic rearrangement of atoms, while the cavity-enhanced lattices enable the creation of large numbers of deep optical potentials that allow for rapid low-loss imaging of atoms. We apply this protocol to demonstrate deterministic filling (99% per-site occupancy) of 1225-site arrays. Because the reservoir is repeatedly filled with fresh atoms, the array can be maintained in a filled state indefinitely. We anticipate that this protocol will be compatible with mid-circuit reloading, which will be a key capability for running large-scale error-corrected quantum computations whose durations exceed the lifetime of a single atom in the system.
@@ I Introduction
Individually controlled neutral atoms provide a promising platform for quantum information processing and simulation, and expanding the size and control of these systems represent key challenges in the ongoing push to access regimes beyond the capabilities of classical simulation. Tweezer arrays and optical lattices have emerged as core technologies for optically trapping and manipulating cold atoms, each with complementary capabilities. Optical lattices provide a well-defined potential landscape with features on the sub-wavelength scale, which can enable tight confinement for imaging and a clean optical potential for simulations involving itinerant atoms [1]. Optical tweezers provide a means of selectively manipulating individual atoms [2]. Notably, tweezer arrays have gained prominence for deterministic assembly of atomic arrays [3, 4, 5] with up to hundreds of atoms [6, 7] and inter-atomic spacing suitable for enabling interactions between atoms through excitation to Rydberg states [8, 9, 10, 6, 7, 11, 12]. Recently, optical tweezers and lattices have been combined to create programmable initial conditions in a Hubbard-regime system [13, 14], and to enable enhanced scaling of atom number for metrology [15] and quantum simulation[16].
The merits of optical lattices can be further enhanced through the use of optical buildup cavities. Optical cavities enhance the lattice depth by enabling the constructive interference of many overlapped reflections of laser light, in turn enabling the creation of arrays of many deep traps [17, 18].
Figure 1: Conceptual diagram of repeated loading sequence. A “reservoir” optical tweezer array (smaller left grid) is repeatedly filled with ${}^{171}$b atoms transported from a spatially separated magneto-optical trap (MOT), and ultimately transferred into a “target” tweezer array (larger right grid) using a “rearrangement” tweezer. The rearrangement moves required for the transfer are determined from low-loss images obtained by transferring atoms from the tweezers into a cavity-enhanced optical lattice and performing site-resolved fluorescence detection. The cavity-enhanced lattice allows for the scalable generation of large numbers of deep traps. The call-out shows example images prior to the 20th cycle of rearrangement, and after the 70th cycle, with the final reservoir reloading step omitted. This protocol allows us to load over 1200 atoms into 1225 target sites.
@@
In this work, we combine the capabilities of tweezers and cavity-enhanced optical lattices to demonstrate an iterative approach to creating large arrays of individually-controlled atoms. Typically, tweezer rearrangement is performed by stochastically loading up to a single atom into each trap within an array [19], imaging the atoms to determine trap occupancy, and then rearranging atoms within the array to create a deterministically occupied sub-array [4, 5]. Crucially, the number of atoms contained in the final array with this approach is no greater than the number initially loaded. Further, because the initial loading is stochastic, the number of sites in the array must generally be substantially larger than desired final sub-array (though under certain conditions, near-deterministic loading can be achieved [20, 21], and several initial arrays can be used to increase the number of available atoms [22]). Recently, repeated loading of a “buffer” array from an optical dipole trap “reservoir” has demonstrated that one can decouple the filling of a six-site target array from a single loading of a cold reservoir [23]. In this work, we extend this concept to repeated loading of a reservoir array, from which we create a deterministically filled target array (typically 99% occupancy) of over 1200 ${}^{171}$b atoms in 1225 sites. This is made possible by combining optical tweezer arrays with a cavity-enhanced optical lattice to provide both microscopic control and the large number of deep traps required for rapid, high-fidelity, low-loss imaging of large numbers of atoms.
A single cycle of loading proceeds as follows (See figure 1): Atoms are collected and cooled in a magneto-optical trap 30 cm below a “science region” where we create a cavity-enhanced optical lattice, as well as tweezer arrays projected through a high-numerical-aperture imaging. The atoms are then transported into the science region, where they are loaded into a set of tweezers that form the reservoir. The reservoir sits adjacent to the target tweezer array, which on all loading cycles but the first is already partially occupied with atoms loaded on previous cycles. The atoms from both reservoir and target tweezer arrays are then transferred into the trapping potential of the cavity-enhanced optical lattice, where we perform site-resolved nondestructive fluorescence imaging to determine which sites of each tweezer array were occupied. We then transfer the atoms back from the cavity lattice to the tweezers, and use a single separate “rearrangement” tweezer to move atoms from the reservoir into empty target tweezers. This defines one loading cycle. Over the course of multiple cycles, the number of atoms in the target array increases until the target array is filled. At this point, further operations may be performed on the atoms, and subsequent loading cycles are applied to maintain a filled array.
@@
Because we reload the reservoir on each loading cycle from a fresh magneto-optical trap, we can continue loading the target array indefinitely. In the near-term, this allows for relatively high data-rates for quantum simulation and computation with large system sizes, and could also be of benefit for optical clocks for high statistical precision with low dead-time [24]. Ultimately, the ability to reload new atoms while maintaining both the presence and coherence of existing atoms will be a key capability for performing error-corrected quantum computations, where execution of an algorithm may take much longer than the lifetime of any given atom in the system. When combined with site-selective hiding [25, 26] and mid-circuit rearrangement techniques already demonstrated in a similar system [26], we anticipate that this protocol will be compatible with mid-circuit reloading of atoms into the array. As a means of maintaining a fully filled array, our approach represents an alternative to the interleaved use of two atomic species [27, 28]. Unlike the two-species approach, the method presented here does not require simultaneous replacement of the entire array, and so may be expected to ease requirements for scaling to large arrays, and would not require inter-species gates [29] in order to replenish the array during a quantum computation.
II Trapping and imaging atoms in a 3D cavity enhanced lattice
Figure 2: Cavity-enhanced optical lattices for rapid low-loss imaging. (a) The intersection of two cavity modes provide three-dimensional confinement for atoms within the field of view of our high-numerical-aperture imaging system. Confinement in the $x$ and $y$ directions is provided by a self-intersecting standing-wave cavity, while confinement in the $z$ direction is provided by a separate self-intersecting running-wave cavity mode. Imaging light (not shown for clarity) is incident along the mode of the running-wave lattice. (b) Observed counts during a 7 ms long image of single atoms within the optical lattice, showing well-resolved peaks for occupied and unoccupied sites. For these conditions, per-image survival is 99.8(1)%. (c) Lattice homogeneity across the target array as characterized by light shifts on the ${}^{1}$${}_{0}$ $m_{f}=1/2$ to ${}^{3}$${}_{1}$, $F=3/2$ $m_{f}=-1/2$ transition. Atoms within 13.4 $\mu$m squares are averaged together for statistics and display clarity. (d, e) Light shift contributions from each cavity averaged over rows (d) and columns (e). Red and orange markers and lines represent the measured and predicted profiles of the XY (red) and Z (orange) cavities, with the prediction centered on the array.
Rapid low-loss imaging of the locations of atoms is essential to our repeated loading protocol. Because atoms are heated by the photon scattering required for imaging, deep optical traps and imaging protocols with low equilibrium temperatures are required to prevent atom loss. Optical tweezers represent a common method for achieving deep optical traps – the tight foci can provide high intensities for moderate powers. However, the total power requirement for a large array of tweezers presents scaling limitations due to limited available laser power. Cavity-enhanced optical lattices offer an alternative path to achieving many deep traps. While the laser intensity in a lattice is spread over a larger cross-sectional area than in optical tweezers both because it spans the regions between utilized sites, and because the Gaussian mode profile must be large compared to the size of the array to achieve uniform trapping, interference of many reflected paths in a cavity-supported lattice can enable more power-efficient generation of traps.
We employ two intersecting optical cavities to provide tight confinement in three dimensions, as shown in figure 2a. The first cavity – which we call the XY cavity – is a retro-reflected four-mirror cavity whose mode intersects itself at the location of the atoms. The polarization is perpendicular to the plane of propagation, providing interference between the two crossed portions of the cavity mode [30]. Compared to using two crossed non-interfering lattices [31, 18], and assuming equal cavity finesse, this configuration provides an eight-fold enhancement in trap depth (defined as the potential barrier between adjacent lattice sites) per unit of total laser power. The XY cavity provides tight confinement in the directions perpendicular to the high numerical aperture optical axis used for the tweezers and imaging. The second cavity, which we call the Z cavity, is a four-mirror running wave cavity whose mode intersects itself at the location of the atoms [32] at 15 degree angles from the XY plane. This provides confinement along the tweezer axis. The polarization of the Z cavity is also oriented perpendicular to its intersection plane, enabling complete interference of the crossing modes. Each cavity has two highly reflecting mirrors, and two matched partially transmitting mirrors. The finesse of the XY and Z cavities are 2900(100) and 3000(100) respectively, and the design mode waists at the location of the atoms are 268 $\mu$m and 183 $\mu$m respectively. For imaging, we typically operate with XY (Z) trap frequencies near 160 kHz (50 kHz), corresponding to 330 $\mu$K (260 $\mu$K) deep traps.
We image atoms using the the narrow-linewidth (180 kHz) ${}^{1}$${}_{0}$ to ${}^{3}$${}_{1}$ transition, which simultaneously provides cooling and scattering of photons for detection [33, 34, 26]. We use a single beam, detuned several hundred kHz from the $F=3/2$, $m_{f}=3/2$ state, with projection onto both the X and Z directions. For applications like array assembly where we wish to determine the presence of an atom but not the state of its nuclear spin (the qubit we use for quantum information applications), we simultaneously apply optical pumping on the ${}^{1}$${}_{0}$ $m_{f}=-1/2$ to ${}^{3}$${}_{1}$ $F=1/2$, $m_{f}=1/2$ transition with light incident along the magnetic field ($x$ direction). At our benchmark imaging parameters, we collect 50 photons in 7 ms, and distinguish occupied from unoccupied sites with a fidelity of approximately 99.95%, as determined from the overlap error in a double-Gaussian fit to the bimodal distribution of counts. Atom loss during such an image is $2(1)\times 10^{-3}$. Spin-selective imaging requires us to omit the optical pumping, and rely on the combination of frequency and polarization selectivity within the ${}^{3}$${}_{1}$ manifold to create a large imbalance of scattering between the two nuclear spin states, as demonstrated in optical tweezers in refs. [34, 26]. The omission of the optical pumping does not modify the photon scattering rate appreciably, but does introduce a finite spin flip probability of of $4(1)\times 10^{-3}$ for our typical imaging parameters, measured as an additional apparent loss in repeated images of ${}^{1}$${}_{0}$ $m_{f}=1/2$. These spin flips are likely due to the fact that the lattice polarization is perpendicular to the magnetic field, and so can induce spin-changing transitions within the ${}^{3}$${}_{1}$ manifold [34].
In order to ensure uniform detuning of the imaging transition across the array and between different motional states, we operate with a trapping wavelength that has equal polarizability for the ground and excited states of the imaging transition – a so-called magic wavelength. For the ${}^{1}$${}_{0}$ $m_{f}=1/2$ to ${}^{3}$${}_{1}$ $m_{f}=3/2$ transition, we find a magic condition near 783.8 nm.
We characterize the homogeneity of the optical lattice potential by probing the ${}^{1}$${}_{0}$ $m_{f}=1/2$ to ${}^{3}$${}_{1}$ $F=3/2$, $m_{f}=-1/2$ transition, for which the excited state has approximately 40% higher polarizability than the ground state. We infer the trap depth from the transition light shifts, measured using optical pumping from the ${}^{1}$${}_{0}$ $m_{f}=1/2$ state, followed by spin-selective imaging (fig. 2c). By varying the power in each cavity, we can extract their independent contributions, displayed averaged over rows and columns in figure  2d,e. Over our 115 $\mu$m square array, peak deviations are within 20% for the Z lattice and 10% for the XY lattice. Because we operate with a magic wavelength condition for critical operations, these deviations do not present a major limitation.
III Repeated loading to build up a large array
@@
Figure 3: (a) Timing diagram for our continuous loading sequence. See main text for a detailed description of the procedure. The beam paths referenced in the call-out are illustrated in figure 4. (b) Number of atoms and fill fraction as a function of loading iteration number during initial loading (left panel) and while maintaining a filled array (right panel). Black points represent atom number inferred from images preceding the rearrangement step. Red points represent atom number inferred from additional images inserted after rearrangement, representing atoms available for computation.
In this section, we describe our continuous loading protocol in detail, and characterize its performance for loading large arrays.
Our loading cycle (illustrated in figure 3a) begins with collecting atoms from a pre-cooled atomic beam into a magneto-optical trap (MOT) using a “core-shell” configuration [35]. In this MOT, each beam consists of a ring of 399 nm light near resonance with the broad linewidth ${}^{1}$${}_{0}$ to ${}^{1}$${}_{1}$ transition surrounding a central region with only 556 nm light near resonance with the narrow linewidth ${}^{1}$${}_{0}$ to ${}^{3}$${}_{1}$ transition. The shell is primarily responsible for capturing atoms from the incident beam, while the low Doppler temperature of the narrow transition used in the core enables cooling to several 10 $\mu$K. Because atoms do not scatter light from the broad transition except during initial capture, the core-shell design poses less risk of generating scattered photons that may be absorbed by atoms in the target array – the total scattering rate on the narrow transition is much lower, and the large magnetic field present in the science region dramatically suppresses reabsorption of photons via the narrow-line transition. In our operating condition, the 399 nm photon scattering rate is about 2 orders of magnitude lower than the broad-line stage of a sequential MOT. We find that the core-shell MOT configuration increases the MOT loading rate by roughly a factor of two, compared to a more standard sequential MOT configuration in a static magnetic field.
From the core-shell MOT, atoms are loaded into a vertically oriented standing-wave lattice formed by counterpropagating 532 nm laser beams focused at the location of the atoms. Loading is achieved by overlapping the lattice with the center of the MOT, followed by reducing the detuning and intensity of the 556 nm light. We then extinguish the MOT light and transport the atoms vertically by synchronously applying a frequency offset to the upward-going and downward-going lattice beams, and translating their foci [26]. This brings the atoms 30 cm vertically into the vicinity of the target array, at which point we transfer the atoms into a reservoir array formed by 483 nm tweezers (another magic wavelength for the ${}^{1}$${}_{0}$ to ${}^{3}$${}_{1}$ transition [26]). To avoid disturbing atoms already loaded into the target array, this transfer takes place at a location displaced 170 $\mu$m horizontally from the target array. No dissipation is required to transfer atoms from the transport lattice to the tweezers – loading is achieved by simply increasing the intensity of the tweezers while the lattice is overlapped and then decreasing the intensity of the lattice. Once loaded, the reservoir array is translated to be directly adjacent to the target array by changing the angle of a galvo mirror that is imaged onto the entrance pupil of the objective.
The target array consists of a rectangular array of tweezer spots formed with light at 459.5960(5) nm wavelength, a magic wavelength for the ${}^{1}$${}_{0}$ to ${}^{3}$${}_{0}$ clock transition [26, 36], and a set of tweezer spots formed by light at 423.31 nm, which is near resonance with a transition between ${}^{3}$${}_{1}$ to a higher-lying 6S8S ${}^{3}$${}_{1}$ state. Both wavelengths provide large light-shifts to atoms in ${}^{3}$${}_{1}$, which prevents unwanted scattering from atoms loaded in these tweezer arrays, and can be used interchangeably for this work. For this work, we operate with a 35x35-site target array with 3.3 $\mu$m spacing, comprised of two separate rectangular arrays each using one of the two previously mentioned wavelengths. We have also operated with overlapped full-size arrays to achieve sufficient trap depth, and obtained similar results. However, the overlapped condition is more sensitive to alignments, and so we use spatially separated arrays for all results presented here.
While translating the reservoir array adjacent to the target array, both arrays are illuminated with light resonant with the ${}^{1}$${}_{0}$ to ${}^{3}$${}_{1}$ $F=3/2$, $m_{f}=3/2$ transition. This induces light-assisted collisions between atoms in the reservoir tweezers, resulting in either zero or one atom in each reservoir tweezer with approximately equal probability [19]. Importantly, the tweezer light illuminating the atoms in the target array prevents them from scattering the light used to induce collisions, which might otherwise cause loss of existing atoms in the target array. Next, atoms from both tweezer arrays are transferred into the optical lattice and imaged, as described above. The imaging leaves the atoms at a temperature of 20 $\mu$K, which is too warm to efficiently transfer into the relatively shallow tweezers of the target array. Following each image, we apply Doppler and Raman sideband cooling to reach an average of $\sim$ 0.1 motional quanta in each direction (Appendix A). Once cooled, the atoms are transferred back into the tweezer arrays, and a single optical tweezer derived from the same laser as the reservoir light is used to transfer atoms from filled sites of the reservoir array into empty sites of the target array. This completes a single loading cycle.
In order to shorten the time required to build up an atomic array, and to limit the losses associated with background gas collisions, we load the MOT at the same time as performing operations in the science chamber. This is made possible by the spatial separation between MOT and science regions, and by the fact that we operate with static magnetic fields: we can simultaneously operate with the magnetic field gradients required by the MOT and the large 500 Gauss homogeneous bias field that is desirable for imaging and other operations required in the science regions. The low scattering rate of 399 nm photons, combined with the large magnetic field and spatial separation between the MOT and science regions prevents light scattered in the MOT from being reabsorbed by atoms in the science region, which could cause heating or decoherence.
@@
Figure 3b shows examples of the atom number increasing as a function of load cycles, with 105 sites in the reservoir and 1225 sites in the target array. At first, the atom number increases approximately linearly as roughly 45 atoms are transferred into the target array per cycle. The loading rate decreases slightly as the array fills, as some atoms from the reservoir are used to counteract loss. Once the number of vacancies in the target array drops below the typical number of atoms in the reservoir, an equilibrium is reached with the vacancy fraction set by the per cycle probability of atom loss.

### v3

# Iterative assembly of 171Yb atom arrays with cavity-enhanced optical lattices

@@ Abstract
Assembling and maintaining large arrays of individually addressable atoms is a key requirement for continued scaling of neutral-atom-based quantum computers and simulators. In this work, we demonstrate a new paradigm for assembly of atomic arrays, based on a synergistic combination of optical tweezers and cavity-enhanced optical lattices, and the incremental filling of a target array from a repetitively filled reservoir. In this protocol, the tweezers provide microscopic rearrangement of atoms, while the cavity-enhanced lattices enable the creation of large numbers of optical traps with sufficient depth for rapid low-loss imaging of atoms. We apply this protocol to demonstrate near-deterministic filling (99% per-site occupancy) of 1225-site arrays of optical traps. Because the reservoir is repeatedly filled with fresh atoms, the array can be maintained in a filled state indefinitely. We anticipate that this protocol will be compatible with mid-circuit reloading of atoms into a quantum processor, which will be a key capability for running large-scale error-corrected quantum computations whose durations exceed the lifetime of a single atom in the system.
@@ I Introduction
Individually controlled neutral atoms provide a promising platform for quantum information processing and simulation, and expanding the size and control of these systems represent key challenges in the ongoing push to access regimes beyond the capabilities of classical simulation. Tweezer arrays and optical lattices have emerged as core technologies for optically trapping and manipulating cold atoms, each with complementary capabilities. Optical lattices provide a well-defined potential landscape with features on the sub-wavelength scale, which can enable tight confinement for imaging and a low-disorder optical potential for simulations involving itinerant atoms [1]. Optical tweezers provide a means of selectively manipulating individual atoms [2]. Notably, tweezer arrays have gained prominence for deterministic assembly of nearly defect-free atomic ensembles [3, 4, 5] with up to hundreds of atoms [6, 7] and inter-atomic spacing suitable for Rydberg-mediated interactions between atoms [8, 9, 10, 6, 7, 11, 12]. Recently, optical tweezers and lattices have been combined to create programmable initial conditions in a Hubbard-regime system [13, 14], and to enable enhanced scaling of atom number for metrology [15] and quantum simulation[16].
The merits of optical lattices can be further enhanced through the use of optical buildup cavities. Optical cavities enhance the lattice depth by enabling the constructive interference of many overlapped reflections of laser light, in turn enabling the creation of large numbers of optical traps with sufficient depth to enable operations that require tight atomic confinement such as low-loss imaging. [17, 18]. In our case, the use of cavity-enhanced lattices provides a roughly two-order-of-magnitude enhancement in power-efficiency for generating traps for imaging, when compared to optical tweezers (See appendix I). In this work, we combine the flexibility and control of optical tweezers with the scalable generation of deep traps afforded by cavity-enhanced optical lattices to demonstrate an iterative approach to creating large arrays of individually-controlled atoms.
Figure 1: Conceptual illustration of repeated loading sequence. (1) Load Reservoir: The first step is to load an array of “reservoir” tweezer traps with 171Yb atoms. These atoms are collected in a spatially separated magneto-optical trap (MOT) (i), transported to the reservoir sites using a moving optical lattice (ii), and then transferred into the reservoir sites (iii). During these steps, atoms loaded on previous iterations of the loading sequence are held in tweezers that form “target” sites. (2) Image Atoms: The second step is to determine the occupancy of reservoir and target sites. This is done by transferring the atoms from the tweezers into a cavity-enhanced optical lattice (i), performing fluorescence imaging within the lattice, as shown after 20 and 70 iterations of loading (ii), and then transferring the atoms back into tweezers. The cavity-enhanced lattice (described in section III) allows for the efficient generation of large numbers of traps for imaging (see appendix I). (3) Load Target Array: The third step is to fill empty target sites with atoms from the reservoir. The preceding image determines which sites are occupied, and a single mobile tweezer drags atoms from occupied reservoir sites to empty target sites. This cycle can be repeated until the target array is filled.
@@
Typically, optical tweezers can be used to assemble defect-free arrays of atoms by stochastically loading up to a single atom into each of a set of tweezers [19], imaging the atoms to determine trap occupancy, and then rearranging atoms within the traps to create a deterministically occupied sub-array [4, 5]. Crucially, the number of atoms contained in the final ensemble with this approach is no greater than the number initially loaded. Further, due to the initial stochastic loading, many extra sites are required to load a sufficient number of atoms to fill the target sub-array, especially when accounting for loss that occurs during the rearrangement process [20] (though under certain conditions, near-deterministic loading can be achieved [21, 22], and several initial arrays can be used to increase the number of available atoms [23]). Recently, repeated loading of a “buffer” array from an optical dipole trap “reservoir” has demonstrated that one can decouple the filling of a six-site target array from a single loading of a cold reservoir [24]. In this work, we extend this concept to repeated loading of a reservoir, from which we create a near-deterministically filled target array (typically 99% occupancy) of over 1200 171Yb atoms in 1225 sites.
Our approach allows us to maintain the filling of our atomic array for an arbitrary duration by replacing atoms that are lost. In the near-term, this allows for relatively high data-rates for quantum simulation and computation with large system sizes, and could also be of benefit for optical clocks for high statistical precision with low dead-time [25].
@@
Ultimately, the ability to reload new atoms while maintaining both the presence and coherence of existing atoms (“mid-circuit reloading”) will be a key capability for performing error-corrected quantum computations, where execution of an algorithm may take much longer than the lifetime of any given atom in the system. When combined with site-selective hiding [26, 27] and mid-circuit rearrangement techniques already demonstrated in a similar system [27], we anticipate that this protocol will be compatible with mid-circuit reloading of atoms.
As a means of maintaining a fully filled array, our approach represents an alternative to the interleaved use of two atomic species [28, 29]. Unlike the two-species approach, the method presented here does not require simultaneous replacement of the entire array, and so may be expected to ease requirements for scaling to large arrays, and would not require inter-species gates [30] in order to replenish the array during a quantum computation.
II Repeated loading to build up a large array
In this section, we provide a high-level description of our experimental apparatus and iterative loading protocol, with further details provided in the appendices. A description of our cavity-enhanced optical lattices and imaging of atoms within the lattices can be found in section III. The iterative protocol can be broken into three main phases, which are illustrated in figure 1 and repeated until the array is filled: 1) Loading atoms into the reservoir tweezers, 2) Imaging the atoms by transferring atoms into the optical lattices and back, and 3) Filling empty sites in a target tweezer array with atoms from the reservoir.
This sequence is enabled by an apparatus that is divided into two regions: a “science” region and a “MOT” (magneto-optical trap) region, which are separated by 30 cm vertically. Atoms are repeatedly collected in a MOT within the MOT region and transported to the science region using an optical lattice formed from counterpropagating 532 nm beams. In the science region, the atoms are accumulated in optical traps. Importantly, each region has a distinct and static magnetic field environment: a magnetic-field gradient in the MOT region, and a large uniform field in the science region (500 Gauss, in the plane of the atom array – the $x$ direction in figure 3a). The static fields and spatial separation allow us to load atoms into the MOT without causing heating or loss of atoms in the science region. This means that the MOT can be loaded while performing other operations in the science region [27]. To further reduce the potential for scattered light from atoms in the MOT region being absorbed by atoms in the science region, we operate the MOT in a “core-shell” configuration [31] which has a dramatically lower scattering rate than a conventional sequential MOT (see appendix A). Details on the optical lattice used to transport atoms to the science region can be found in appendix B.
Within the science region, several sets of optical traps facilitate the accumulation of atoms into a desired configuration. These traps are either intensity maxima of optical tweezers, formed by tightly focusing laser beams through a high-numerical-aperture (NA = 0.65, field of view = 0.5 mm) microscope objective, or of co-located cavity-enhanced optical lattices, described in section III. First, atoms are transferred from the transport lattice into a set of reservoir tweezers, where they are cooled and undergo light-assisted collisions (LACs) that ensure stochastic loading of zero or one atom per trap [19]. In addition to the reservoir tweezers, the science region contains a square grid of 1225 “target array” tweezers, where we accumulate atoms. Importantly, we use a wavelength for the reservoir tweezers (483 nm) that does not shift the frequency of the narrow-linewidth 1S0 $m_{F}=1/2$ to 3P1, $F=3/2$ $m_{F}=3/2$ transition used for cooling and LACs (referred to as a “magic wavelength”). In contrast, the target array uses light at wavelengths (423 nm and 460 nm) that cause a large shift on the cooling and LAC transition, dramatically reducing the scattering rate and loss of atoms already loaded in the target array during cooling and LAC. Further details on the reservoir and target array traps can be found in appendices B and  C.
After loading the reservoir, the next step is to transfer the atoms into vacant target tweezers. This requires identifying both which sites in the reservoir are occupied, and which sites in the target are vacant. To do this, atoms from both tweezer arrays are transferred to the cavity-enhanced optical lattice, where they are imaged with low loss (described in section III), and then transferred back into the tweezer arrays. Based on the information gained from the image about which tweezer sites were unoccupied (either because they had yet to be filled or because a previously loaded atom had been lost), atoms are then moved from the reservoir into unoccupied target sites using a single “rearrangement” optical tweezer (described in appendix D). The cycle is then repeated until the target array is filled. At this point, further operations may be performed on the atoms, and subsequent loading cycles are applied to maintain a filled array.
@@ II Repeated loading to build up a large array
Figure 2: (a) Number of atoms and fill fraction as a function of loading iteration number during initial loading. (b) The same quantities while maintaining a filled array. Black markers represent the atom number and fill fraction determined from images taken before the rearrangement step. Red markers represent images taken after rearrangement on select iterations, which indicates the number of atoms available for computation (see main text for further details).
@@
Figure 2a shows an example of the atom number increasing as a function of load cycles, with 105 sites in the reservoir and 1225 sites in the target array. At first, the atom number increases approximately linearly as roughly 45 atoms are transferred into the target array per cycle. The loading rate decreases slightly as the array fills, as some atoms from the reservoir are used to counteract loss. Once the number of vacancies in the target array drops below the typical number of atoms in the reservoir, an equilibrium is reached with the vacancy fraction set by the per cycle probability of atom loss. Figure 2b shows the fill fraction while maintaining a filled array by continuing to repeat the loading protocol.